### v1

# Curves of genus two with maps of every degree to a fixed elliptic curve

Abstract.
We show that up to isomorphism there are exactly twenty pairs $(C,E)$, where $C$ is a genus-$2$ curve over $\mathbf{C}$}, where $E$ is an elliptic curve over $\mathbf{C}$}, and where for every integer $n>1$ there is a map of degree $n$ from $C$ to $E$.
Key words and phrases:
Hyperelliptic curve, elliptic curve, abelian surface, split Jacobian, quadratic form
1991 Mathematics Subject Classification:
Primary 14H45; Secondary 11G05, 11G10, 11G15, 11G30
1. Introduction
Curves of genus two that have nonconstant maps to elliptic curves have been studied for nearly $200$ years, beginning with work of Legendre in 1828. Below, we will briefly review some of the work of the early researchers in the field — Legendre, Jacobi, Weierstrass, Kowalevski, Poincaré, Picard, Goursat, Briochi, and others — but for now we will simply note that the problem we consider in this paper is one that could be understood by these authors, with just a little tweaking of the terminology. Namely, we address the question of whether there exists a genus-$2$ curve $C$ over the complex numbers $\mathbf{C}$}, and an elliptic curve $E$ over $\mathbf{C}$}, such that for every $n>1$ there exists a degree-$n$ morphism from $C$ to $E$. (“Is there a hyperelliptic integral that can be reduced, via transformations of every degree $n>1$, to expressions involving the same elliptic integral?”)
Perhaps surprisingly, the answer is yes.
Theorem 1.
Up to isomorphism, there are exactly twenty pairs $(C,E)$ such that
  1. (1)
    $C$ is a curve of genus $2$ over the complex numbers $\mathbf{C}$};
  2. (2)
    $E$ is an elliptic curve over $\mathbf{C}$}; and
  3. (3)
    for every $n>1$ there is a map of degree $n$ from $C$ to $E$.
Suppose $(C,E)$ is one of these twenty pairs. If we choose a base point $P$ on $C$, then the set of maps from $C$ to $E$ that take $P$ to the origin of $E$ is a free $\mathbf{Z}$}-module of rank $4$, and the degree function is a quadratic form on this module. The twenty pairs give rise to only four different quadratic forms on ${\mathbf{Z}}^{4}$, up to isomorphism. These quadratic forms are
$q_{1}$ $=2w^{2}+3x^{2}+3y^{2}+4z^{2}+2xy$
$q_{2}$ $=2w^{2}+2x^{2}+3y^{2}+3z^{2}+2wz+2xy$
$q_{3}$ $=2w^{2}+3x^{2}+3y^{2}+4z^{2}+2wx+2wy+2xz+2yz$
$q_{4}$ $=2w^{2}+3x^{2}+4y^{2}+6z^{2}-2wx+2wz+2xy+4yz,$
and so in the course of proving Theorem 1 we will need the following result.
Proposition 2.
Each of the quaternary quadratic forms $q_{1}$, $q_{2}$, $q_{3}$, $q_{4}$ represents every integer greater than $1$.
1-11/2-1/201/21/211ii
Figure 1. A strict fundamental domain ${\mathcal{F}}_{1}$ for $\Gamma(1)$
Let ${\mathcal{F}}_{1}$ be the strict fundamental domain for $\Gamma(1)$ depicted in Figure 1. Suppose $(C,E)$ is one of the twenty pairs from Theorem 1, and let $\tau$ be the element of ${\mathcal{F}}_{1}$ that corresponds to $E$. We will show that $E$ has complex multiplication, so that $\tau$ is an element of an imaginary quadratic field. We will also show that the curve $C$ has a period matrix of the form
$\begin{pmatrix}1&0&\tau/2&1/2\\ 0&1&1/2&\sigma/2\end{pmatrix}$
where $\tau$ is as above and where $\sigma$ lies in the strict fundamental domain ${\mathcal{F}}_{2}$ for $\Gamma(2)$ depicted in Figure 2. (That there is a period matrix of this form, for any genus-$2$ curve with a map of degree $2$ to an elliptic curve, is essentially a result of Picard [30, 31].) Table 1 gives the value of $\tau$ and $\sigma$ for the each of the twenty pairs, along with the discriminants $\Delta_{E}$ and $\Delta_{F}$ of the endomorphism rings of $E$ and $F$, and the quadratic form associated to the pair $(C,E)$. Pairs $(C,E)$ that have the same values of $\Delta_{E}$ and $\Delta_{F}$ can be obtained from one another by Galois conjugation.
[…]
Figure 2. A strict fundamental domain ${\mathcal{F}}_{2}$ for $\Gamma(2)$, whose closure is tiled with images of the closure of the strict fundamental domain ${\mathcal{F}}_{1}$. The tiles are labeled by the Möbius transformation that takes ${\mathcal{F}}_{1}$ to the given tile. Note that $(3+\sqrt{-3})/2$ and $(3+\sqrt{-3})/6$ are not included in ${\mathcal{F}}_{2}$.
No. $\Delta_{E}$ $\Delta_{F}$ $\tau$ $\sigma$ Form
1. 4-4 100-100 $\sqrt{-1}$ $5\sqrt{-1}$ $q_{2}$
2. $(12+5\sqrt{-1})/13$
3. 8-8 32-32 $\sqrt{-2}$ $\sqrt{-2}/4$ $q_{1}$
4. $(4+\sqrt{-2})/4$
5. $(1+\sqrt{-2})/2$
6. $(2+\sqrt{-2})/4$
7. 72-72 $(6+\sqrt{-2})/6$ $q_{3}$
8. $(2+3\sqrt{-2})/2$
9. 12-12 3-3 $\sqrt{-3}$ $(-1+\sqrt{-3})/2$ $q_{2}$
10. $(1+\sqrt{-3})/2$
11. 16-16 4-4 $2\sqrt{-1}$ $\sqrt{-1}$ $q_{1}$
12. $1+\sqrt{-1}$
13. 20-20 20-20 $\sqrt{-5}$ $\sqrt{-5}$ $q_{2}$
14. $(1+\sqrt{-5})/2$ $(1+\sqrt{-5})/2$
15. 24-24 24-24 $\sqrt{-6}$ $(2+\sqrt{-6})/2$ $q_{3}$
16. $\sqrt{-6}/2$ $(6+\sqrt{-6})/7$
17. 36-36 36-36 $3\sqrt{-1}$ $(6+3\sqrt{-1})/5$ $q_{4}$
18. $(4+3\sqrt{-1})/5$
19. $(1+3\sqrt{-1})/2$ $1+3\sqrt{-1}$
20. $(3+\sqrt{-1})/3$
Table 1. Data for the twenty pairs $(C,E)$ from Theorem 1
The structure of this paper follows that of the proof of Theorem 1. In Section 2 we recall some facts about genus-$2$ curves with degree-$2$ maps to elliptic curves. In particular, the following proposition is fundamental to our proof.
Proposition 3.
Suppose $C$ is a genus-$2$ curve with a degree-$2$ map $\varphi$ to an elliptic curve $E$. Then there is a unique elliptic curve $F$, a degree-$2$ map $\chi\colon C\to F$, and an isomorphism $\psi\colon E[2]\to F[2]$ such that the kernel of $\varphi^{*}\times\chi^{*}\colon E\times F\to\operatorname{Jac}C$ is the graph of $\psi$ and such that the following diagram commutes:
[…]
The pair $(\chi,\psi)$ is unique up to composition with automorphisms of $F$. Conversely, given two elliptic curves $E$ and $F$ and an isomorphism $\psi\colon E[2]\to F[2]$, there is a genus-$2$ curve $C$ and a degree-$2$ map $C\to E$ that gives rise to $F$ and $\psi$ as above, unless $\psi$ is the restriction to $E[2]$ of an isomorphism $E\to F$, in which case there is no such curve $C$.
Corollary 4.
Let notation be as in Proposition 3, and let $\omega$ be a nonconstant map from $C$ to $E$. Let $\alpha$ be the isogeny $\varphi_{*}\omega^{*}\colon E\to E$ and let $\beta$ be the morphism $\chi_{*}\omega^{*}\colon E\to F$ (which is an isogeny if it is nonzero). Then $\deg\omega=(\deg\alpha+\deg\beta)/2$, and for every $P\in E[2]$ we have $\beta(P)=\psi(\alpha(P))$.
There is a converse statement, as well.
Corollary 5.
Let notation be as in Proposition 3, and suppose $\alpha\colon E\to E$ and $\beta\colon E\to F$ are morphisms such that for every $P\in E[2]$, we have $\beta(P)=\psi(\alpha(P))$. Then there is a nonconstant map $\omega\colon C\to E$ with $\deg\omega=(\deg\alpha+\deg\beta)/2$ such that $\alpha=\varphi_{*}\omega^{*}$ and $\beta=\chi_{*}\omega^{*}$.
Suppose $C$ is a genus-$2$ curve that has maps of every degree $n>1$ to an elliptic curve $E$. In Section 3 we apply Corollary 4 to the degree-$3$ and degree-$4$ maps from $C$ to $E$, and use the resulting information to deduce restrictions on the endomorphism ring of $E$ and on the relationship between $E$ and $F$. In particular, we prove the following proposition.
Proposition 6.
Suppose $C$ is a genus-$2$ curve that has maps of degree $2$, 33, and $4$ to an elliptic curve $E$. Let $F$ be the curve associated as in Proposition 3 to a degree-$2$ map from $C$ to $E$. Then for one of the possibilities for $p$ and $\Delta$ listed below, the endomorphism ring of $E$ has discriminant $\Delta$, and there is a cyclic isogeny from $E$ to $F$ of degree $p$.
  1. (1)
    $p=1$ and $\Delta\in\{3$, $4$, 77, $11$, $12$, $16$, 1919, 2020, $24$, $27$, 28}.28\}.
  2. (2)
    $p=2$ and $\Delta\in\{4$, $7$, 88, $12$, $15$, $16$, 2020, $23$, $24$, $31$, $36$, $39$, 40}.40\}.
  3. (3)
    $p=3$ and $\Delta\in\{3$, $4$, 88, $11$, $12$, $16$, 1919, 20}.20\}.
  4. (4)
    $p=5$ and $\Delta\in\{3$, $4$, 77, 88, $11$, $12$, $15$, $16$, 1919, $31$, $35$, $40$, $76$, $91$, $104$, $115$, $124$, $131$, $136$, $139$, 140}140\}.
After Proposition 6, we see that only finitely many pairs $(E,F)$ can occur. Suppose $(E,F)$ is one such pair, and suppose $\psi$ is one of the six isomorphisms $E[2]\to F[2]$. We can compute the curve $C$ associated to this data as in the second statement of Proposition 3, if such a $C$ exists. For each such $C$ and $E$, we can compute a $\mathbf{Z}$}-basis for $\operatorname{Hom}(C,E)$, and using Corollary 5 we can compute the positive definite quadratic form given by the degree map. It is then an easy matter to check whether this form represents all integers $n$ with $1<n<32$, which is obviously a necessary condition for the form to represent all integers $n>1$. Our method for doing this is explained in Section 4, and Magma code for carrying out the computation is available on the GitHub repository mentioned in Section 4.
It turns out that each quadratic forms arising in this way that represent all the integers from $2$ to $31$ is equivalent to one of the forms $q_{1}$, $q_{2}$, $q_{3}$, and $q_{4}$ given above, and therefore the $(C,E)$ pairs that we have found satisfy the conditions of Theorem 1. It is then a simple matter to compute the data presented in Table 1, and to see that there are only $20$ such pairs.
In Section 5 we compute models for the curves $C$. We close in Section 6 by proving Proposition 2.
Remark 7.
We note here that there is a result similar to Theorem 1 for fields of positive characteristic, if we restrict our attention to ordinary curves. Namely, if $C$ is an ordinary genus-$2$ curve over an algebraically closed field $K$ of positive characteristic, and if $C$ has maps of every degree $n>1$ to an elliptic curve $E$, then $C$ is the reduction of one of the curves from Theorem 1. This follows from the fact that the category of ordinary abelian varieties over $K$ embeds into the category of abelian varieties over $\mathbf{C}$}, by a result of Deligne [8]. Note, however, that in general not all of the curves in the theorem will have good ordinary reduction, so there will not necessarily be $20$ examples of such curves over a given $K$.
We have not investigated the situation for non-ordinary curves over a field of positive characteristic. The endomorphism ring of a supersingular elliptic curve is a $\mathbf{Z}$}-module of rank $4$, so in some sense it is easier for there to exist maps $\alpha$ and $\beta$ as in Corollary 5 that can produce an $\omega$ of a given degree. For this reason, we expect that over some fields there will be examples of $(C,E)$ pairs that are not reductions of our $20$ curves in characteristic zero.
Remark 8.
Here we give some historical background. As we mentioned at the beginning of this section, the study of genus-$2$ curves with maps to elliptic curves goes back nearly two centuries. In §12 of the third supplement to his Traité des fonctions elliptiques [28], published in 1828, Legendre shows how several “ultra-elliptic” integrals involving expressions of the form $\sqrt{x(1-x^{2})(1-k^{2}x^{2})}$ can be expressed in terms of elliptic integrals. Jacobi, in a postscript to his 1832 review [19, 20] of Legendre’s book, notes that Legendre’s examples can be generalized; rephrased in modern terminology, Jacobi’s observation is that every hyperelliptic curve of the form
$y^{2}=x(x-1)(x-\lambda)(x-\mu)(x-\lambda\mu)$
admits a degree-$2$ map to an elliptic curve. Legendre’s examples come by taking $\lambda=-1$. Later, Königsberger [23] and Picard [30, §9] each proved that every genus-$2$ curve with a degree-$2$ map to an elliptic curve occurs in Jacobi’s family.
The study of genus-$2$ curves with maps to elliptic curves continued, and flourished, in the latter half of the $19$th century, with the focus shifting to the period matrices of such curves and the endomorphism rings of their Jacobians. In an 1874 paper, not published in a journal until 1884, Kowalevski [24] quotes an unpublished result of Weierstrass that describes the period matrices of curves whose associated abelian integrals can be reduced to elliptic integrals; in 1884 Poincaré [32] provided a proof of Weierstrass’s theorem. For the special case of genus-$2$ curves, a better version of Weierstrass’s result was given (independently) by Picard [30], and in 1884 Picard showed that his result can also be deduced directly from that of Weierstrass [31]. At the very end of the $19$th century, Humbert published a series of papers [16, 17, 18] concerning genus-$2$ curves whose Jacobians have endomorphism rings larger than $\mathbf{Z}$}; Humbert’s curves having “singular relations with square invariant” have zero-divisors in their endomorphism rings, and hence have maps to elliptic curves.
Research in these matters has continued to this day. In more modern terminology, one can fix an integer $n>1$ and study the moduli space of triples $(C,E,\varphi)$, where $\varphi\colon C\to E$ is a map of degree $n$ from a curve of genus $2$ to an elliptic curve. (Usually one demands in addition that the map be minimal, meaning that it does not factor through an isogeny $F\to E$ of degree greater than $1$.) Some work concerns the general case (see for example [11, 21, 22]), but there is also interest in considering specific small values of $n$ and constructing more or less explicit models of the corresponding moduli space, perhaps also giving equations for the triples $(C,E,\varphi)$ themselves.
For $n=2$, Jacobi’s previously-cited work gives such equations over algebraically closed fields; in [15], the authors analyze the situation over non-algebraically closed fields. For the case $n=3$, there are works spanning 141 years, including [2, 3, 4, 5, 12, 13, 14, 25, 33]. The case $n=4$ is considered in older [1] and more recent [7] research, and there is work on the case $n=5$ as well [29]. The paper [26] considers all $n$ up to $11$, but is more focused on models for the moduli space itself rather than on the triples $(C,E,\varphi)$, partly because the known models for $C$ become quite complicated even for $n=4$.
2. Consequences of the existence of a degree-$2$ map
In this section we prove Proposition 3 and its corollaries.
Proof of Proposition 3.
Suppose $C$ is a genus-$2$ curve over $\mathbf{C}$} with a degree-$2$ map $\varphi$ to an elliptic curve $E$. Then the special case $N=2$ of [21, Theorem 1.5] shows that there is another elliptic curve $F$ and an isomorphism $\psi\colon E[2]\to F[2]$ such that the Jacobian of $C$ is isomorphic to the quotient of $E\times F$ by the graph of $\psi$, and such that there is a degree-$2$ map $\chi\colon C\to F$.
Furthermore, if we let $G\subset(E\times F)[2]$ be the graph of $\psi$, then there is an isogeny $\Phi\colon E\times F\to\operatorname{Jac}C$ with kernel $G$, and we have a diagram
[…]
We can then extend this diagram so that the compositions of the horizontal arrows on the top line and on the bottom line are the multiplication-by-$2$ maps:
[…]
The right half of this diagram is nothing other than diagram (1), which is what we want to show exists. The uniqueness of the pair $(\chi,\psi)$ up to automorphisms of $F$ is part of [21, Theorem 1.5], and the converse follows from this as well. ∎
Remark 9.
We note that we can give a period matrix for the Jacobian of $C$ in terms of the period matrices for $E$ and $F$ and the isomorphism $\psi$, as follows. First, $E$ has a period lattice $\Lambda_{E}$ of the form $\langle 1,\tau\rangle$ for a unique $\tau$ in the fundamental domain ${\mathcal{F}}_{1}$, and there is a unique $\sigma$ in the fundamental domain ${\mathcal{F}}_{2}$ such that
  • $\Lambda_{F}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\langle 1,\sigma\rangle$ is a period matrix for $F$, and
  • the isomorphism $\psi\colon E[2]\to F[2]$ sends the $2$-torsion point $1/2\bmod\Lambda_{E}$ of $E({\mathbf{C}})$ to the $2$-torsion point $\sigma/2\bmod\Lambda_{F}$ of $F({\mathbf{C}})$, and the point $\tau/2\bmod\Lambda_{E}$ of $E({\mathbf{C}})$ to the point $1/2\bmod\Lambda_{F}$ of $F({\mathbf{C}})$.
Then we can take
(2) $\Lambda_{C}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\begin{pmatrix}1&0&\tau/2&1/2\\ 0&1&1/2&\sigma/2\end{pmatrix}$
to be a period matrix for the Jacobian of $C$. This is essentially a result of Picard; see [30] and [31].
We also know the sesquilinear form on ${\mathbf{C}}^{2}$ that represents the principal polarization on $\operatorname{Jac}C$, because it is derived from the product polarization on $\Lambda_{E}\times\Lambda_{F}$. Namely, if $\delta$ is any multiple of $\sqrt{-1}$ and we write $\tau=a+b\delta$ and $\sigma=c+d\delta$ for real numbers $a,b,c,d$, then the sesquilinear form applied to elements $(z_{1},z_{2})$ and $(w_{1},w_{2})$ of ${\mathbf{C}}^{2}$ gives the value
(3) $\operatorname{Trace}_{{\mathbf{C}}/{\mathbf{R}}}\Big(\frac{w_{1}\bar{z}_{1}}{b\delta}+\frac{w_{2}\bar{z}_{2}}{d\delta}\Big).$
One can check that the matrix of values of this pairing, applied to pairs of column vectors in the basis for $\Lambda_{C}$ given above, is
$\begin{pmatrix}[r]0&\phantom{-}0&-1&0\\ 0&0&0&-1\\ 1&0&0&0\\ 0&1&0&0\end{pmatrix}$,
so the pairing does indeed give a principal polarization on $\Lambda_{C}$.
Proof of Corollary 4.
Let $\alpha=\varphi_{*}\omega^{*}$ and $\beta=\chi_{*}\omega^{*}$. We can extend diagram (1) as follows:
[…]
Following the bottom edge of the diagram gives us multiplication by $2\deg\omega$ on $E$. The map from $E$ to $E$ we get from following the top edges of the diagram is the sum of the endomorphisms $\widehat{\alpha}\alpha$ and $\widehat{\beta}\beta$ of $E$, where $\widehat{\alpha}$ and $\widehat{\beta}$ are the dual morphisms of $\alpha$ and $\beta$. But $\widehat{\alpha}\alpha$ is multiplication by $\deg\alpha$, and $\widehat{\beta}\beta$ is multiplication by $\deg\beta$, so we see that $2\deg\omega=\deg\alpha+\deg\beta$, as claimed.
Let $P$ be point of order $2$ on $E$. Then the image of $P$ under the map from the lower left of the diagram to the $E\times F$ on the upper right is the pair $(\alpha(P),\beta(P))$, while the image of $P$ in rightmost copy of $\operatorname{Jac}C$ is 0, because the middle map from $\operatorname{Jac}C$ to $\operatorname{Jac}C$ is multiplication by $2$. Therefore, $(\alpha(P),\beta(P))$ lies in the kernel of the isogeny $\varphi^{*}\times\chi^{*}$, which is the graph of $\psi$, and it follows that $\beta(P)=\psi(\alpha(P))$. ∎
Proof of Corollary 5.
Given $\alpha$ and $\beta$ as in the statement of the corollary, consider the following diagram:
[…]
Our goal is to produce a morphism $\omega\colon C\to E$ that will allow us to extend this diagram to diagram (4).
By assumption, we have $\beta(P)=\psi(\alpha(P))$ for every $P\in E[2]$, so the kernel of the map $\widehat{\alpha}+\widehat{\beta}$ from $E\times F$ to $E$ contains the kernel of $\varphi^{*}\times\chi^{*}$. It follows that there is a map $\varpi\colon\operatorname{Jac}C\to E$ that we can use to complete the triangle on the right-hand side of (5). (We note that this map is unique, because $\varphi^{*}\times\chi^{*}$ is an isogeny.)
Choose an Abel–Jacobi embedding of $C$ into its Jacobian, and let $\omega$ be the composition of this embedding with $\varpi$. Then we automatically have $\varpi=\omega_{*}$, and by duality we find that $\omega^{*}\colon E\to\operatorname{Jac}C$ completes the triangle on the left-hand side of (5). This gives us (4), and proves the corollary. ∎
3. Consequences of the existence of maps of degree $3$ and 44
In this section we prove Proposition 6. We begin with a lemma that records some facts about endomorphism rings of elliptic curves with noncylic endomorphisms of small degree.
Lemma 10.
Let $E$ be an elliptic curve over $\mathbf{C}$} that has a cyclic isogeny $\alpha$, and let $\Delta$ be the discriminant of the endomorphism ring of $E$.
  1. (1)
    If $\deg\alpha=2$, then $\Delta\in\{4$, $7$, 8}8\}.
  2. (2)
    If $\deg\alpha=3$, then $\Delta\in\{3$, $8$, $11$, 12}12\}.
  3. (3)
    If $\deg\alpha=4$, then $\Delta\in\{7$, $12$, $15$, 16}16\}.
  4. (4)
    If $\deg\alpha=5$, then $\Delta\in\{4$, $11$, $16$, 1919, 20}20\}.
  5. (5)
    If $\deg\alpha=6$, then $\Delta\in\{8$, $15$, 2020, $23$, 24}24\}.
  6. (6)
    If $\deg\alpha=7$, then $\Delta\in\{3$, $7$, $12$, 1919, $24$, $27$, 28}28\}.
  7. (7)
    If $\deg\alpha=10$, then $\Delta\in\{4$, $15$, $24$, $31$, $36$, $39$, 40}40\}.
  8. (8)
    If $\deg\alpha=35$, then $\Delta\in\{19$, $31$, $35$, $40$, $59$, $76$, $91,104$, $115$, $124$, $131$, $136,$ $139$, 140}140\}.
Proof.
Let $E$ be an elliptic curve whose endomorphism ring is isomorphic to the imaginary quadratic order $\mathcal{O}$} of discriminant $\Delta$. The norms of elements of $\mathcal{O}$} are the integers of the form $x^{2}+\Delta xy+(\Delta^{2}-\Delta)y^{2}/4$, with $x$ and $y$ in $\mathbf{Z}$}. The isogeny corresponding to a given $x$ and $y$ is cyclic if and only if $x$ and $y$ are coprime to one another. Given a norm $n$, it is a simple matter to find the $\Delta$ for which there exist coprime $x$ and $y$ giving an element of norm $n$. We leave the details to the reader. ∎
Proof of Proposition 6.
Suppose $C$ is a genus-$2$ curve over $\mathbf{C}$} that has maps of degree $2$, 33, and $4$ to an elliptic curve $E$. Let $\varphi$ be a degree-$2$ map from $C$ to $E$, and let the elliptic curve $F$, the degree-$2$ map $\chi\colon C\to F$, and the isomorphism $\psi\colon E[2]\to F[2]$ be as in Proposition 3.
By Corollary 4, the existence of the degree-$3$ map from $C$ to $E$ implies that there is an endomorphism $\alpha$ of $E$ and a morphism $\beta\colon E\to F$ such that $\deg\alpha+\deg\beta=6,$ and such that
(6) $\beta(P)=\psi(\alpha(P))\text{\quad for all $P\in E[2]$.}$
In particular, (6) implies that
(7) $(\ker\alpha)[2]=\#(\ker\beta)[2].$
We enumerate the possibilities below. Note that Lemma 10 tells us the possible discriminants of the endomorphism ring of an elliptic curve with a cyclic isogeny of certain degrees, and we use this without further comment in the list below to indicate how each possibility is covered by one of the cases in the statement of the proposition.
  1. 1.
    $\deg\alpha=0$ and $\deg\beta=6$. This cannot happen, because $(\ker\alpha)[2]=4$ while $(\ker\beta)[2]=2$, contradicting (7).
  2. 2.
    $\deg\alpha=1$ and $\deg\beta=5$. This implies that $F$ is $5$-isogenous to $E$. We explore this case further in the discussion below.
  3. 3.
    $\deg\alpha=2$ and $\deg\beta=4$. By (7), we see that $\beta$ must be a cyclic isogeny. More specifically, (6) implies that $\ker\alpha$ is contained in $\ker\beta$, so $\beta$ is the composition of $\alpha$ with a $2$-isogeny from $E$ to $F$. This possibility therefore falls under the case $p=2$ of the statement of the proposition.
  4. 4.
    $\deg\alpha=3$ and $\deg\beta=3$. This falls under the case $p=3$ of the statement of the proposition.
  5. 5.
    $\deg\alpha=4$ and $\deg\beta=2$. We see from (7) that $\alpha$ must be a cyclic isogeny. Therefore this falls under the case $p=2$ of the statement of the proposition.
  6. 6.
    $\deg\alpha=5$ and $\deg\beta=1$. This falls under the case $p=1$ of the statement of the proposition.
  7. 7.
    $\deg\alpha=6$ and $\deg\beta=0$. Equation (7) shows that this case cannot occur.
The only possibility not covered by the conclusion of the proposition is that $E$ is arbitrary and $F$ is $5$-isogenous to $E$. For the rest of the proof we will assume that we are in this case.
Now we consider the consequences of the existence of a degree-$4$ map from $C$ to $E$. Corollary 4 implies that there is an endomorphism $\alpha$ of $E$ and a morphism $\beta\colon E\to F$ such that $\deg\alpha+\deg\beta=8,$ with (6) and (7) holding. We list the possibilities, and again use Lemma 10 without comment to show which cases of the proposition covers them.
  1. 1.
    $\deg\alpha=0$ and $\deg\beta=8$. From (7) we see that $\ker\beta$ must contain $E[2]$, so $\beta$ is the composition of a $2$-isogeny $E\to F$ with the multiplication-by-$2$ map on $E$. We see that $F$ must be $2$-isogenous to $E$. Since $F$ is also $5$-isogenous to $E$, we see that $E$ has an endomorphism of degree $10$. We find that this possibility falls under the case $p=2$ of the proposition.
  2. 2.
    $\deg\alpha=1$ and $\deg\beta=7$. We will discuss this case below.
  3. 3.
    $\deg\alpha=2$ and $\deg\beta=6$. This falls under the case $p=5$ of the statement of the proposition.
  4. 4.
    $\deg\alpha=3$ and $\deg\beta=5$. This falls under the case $p=5$ of the statement of the proposition.
  5. 5.
    $\deg\alpha=4$ and $\deg\beta=4$. If $\alpha$ is cyclic, then this falls under the case $p=5$ of the statement of the proposition. If $\alpha$ is not cyclic, then by (7) neither is $\beta$, which means that $F\cong E$. Therefore, there is an endomorphism of $E$ of degree $5$. This falls under the case $p=1$ of the statement of the proposition.
  6. 6.
    $\deg\alpha=5$ and $\deg\beta=3$. This falls under the case $p=3$ of the statement of the proposition.
  7. 7.
    $\deg\alpha=6$ and $\deg\beta=2$. This falls under the case $p=5$ of the statement of the proposition.
  8. 8.
    $\deg\alpha=7$ and $\deg\beta=1$. This falls under the case $n=7$, $p=1$ of the statement of the proposition.
  9. 9.
    $\deg\alpha=8$ and $\deg\beta=0$. By (7) we see that $\alpha$ cannot be cyclic, so it is the composition of multiplication-by-$2$ with an endomorphism of $E$ of degree $2$. This falls under the case $p=5$ of the statement of the proposition.
This leaves us with one situation unaddressed: when there is both a $5$-isogeny from $E$ to $F$ and a $7$-isogeny from $E$ to $F$, so that $E$ has an endomorphism of degree $35$. We cannot simply use Lemma 10 to deduce that this situation falls under the case $p=5$ of the proposition, because Lemma 10 says that an elliptic curve with endomorphism discriminant 59-59 can have a cyclic isogeny of degree $35$, but 59-59 is not one of the discriminants the proposition includes for the case $p=5$. The proposition does allow all of the other discriminants listed in case (8) of Lemma 10, so all we must do is to show that the discriminant 59-59 cannot occur.
So suppose, in the situation of the proposition, that the elliptic curve $E$ has endomorphism ring with discriminant 59-59 and that $F$ is both $5$-isogenous and $7$-isogenous to $E$. Since there are isogenies from $E$ to $F$ of coprime degrees, $F$ must also have endomorphism ring with discriminant 59-59. We check that then there are morphisms from $E$ to $F$ of degrees 0, 33, 55, and $7$, and no other degrees less than $9$, and that there are endomorphisms of $E$ of degrees 0, 11, and $4$, and no other degrees less than $9$.
Since we are assuming that there is a degree-$3$ map from $C$ to $E$, Corollary 4 says that there is an $\alpha_{3}\in\operatorname{End}E$ and a $\beta_{3}\in\operatorname{Hom}(E,F)$ such that $\deg\alpha_{3}+\deg\beta_{3}=6$ and such that $\beta_{3}(P)=\psi(\alpha_{3}(P))$ for all $P\in E[2]$. The only possibility is that $\deg\alpha_{3}=1$ and $\deg\beta_{3}=5$. Note that $\alpha_{3}=\pm 1$, so in fact $\beta_{3}(P)=\psi(P)$ for all $P\in E[2]$.
Likewise, since there is a degree-$4$ map from $C$ to $E$ there is an $\alpha_{4}\in\operatorname{End}E$ and a $\beta_{4}\in\operatorname{Hom}(E,F)$ such that $\deg\alpha_{4}+\deg\beta_{4}=8$ and such that $\beta_{4}(P)=\psi(\alpha_{4}(P))$ for all $P\in E[2]$. The only possibility is $\deg\alpha_{4}=1$ and $\deg\beta_{4}=7$, so that $\alpha_{4}=\pm 1$ and $\beta_{4}(P)=\psi(P)$ for all $P\in E[2]$.
If we let $\widehat{\beta}_{4}$ be the dual isogeny of $\beta_{4}$, so that $\widehat{\beta}_{4}\beta_{4}=7$, then $P=\widehat{\beta}_{4}(\psi(P))$ for all $P\in E[2]$. Therefore, if we set $\gamma=\widehat{\beta}_{4}\beta_{3}$, then $\gamma$ is a degree-$35$ endomorphism of $E$ that acts trivially on $E[2]$. That implies that $\gamma-1$ kills $E[2]$, so $\gamma-1=2\delta$ for some $\delta\in\operatorname{End}E$.
But we check that the only elements of norm $35$ in $\operatorname{End}E\cong{\mathbf{Z}}[\frac{1+\sqrt{-59}}{2}]$ are $\frac{\pm 9\pm\sqrt{-59}}{2}$, and none of these can be written as $1+2\delta$ for $\delta\in{\mathbf{Z}}[\frac{1+\sqrt{-59}}{2}]$. Therefore, $\operatorname{End}E$ cannot have discriminant 59-59, and our last remaining situation is covered by the case $p=5$ of the proposition. ∎
Remark 11.
We choose to exclude the discriminant 59-59 from the statement of Proposition 6, even though its exclusion complicates the proof, because including it would make one of our later computational steps slightly more awkward.
4. Enumerating possible examples
Proposition 6 gives our first step toward our proof of Theorem 1 by specifying a finite list of possible pairs $(E,F)$ from which to construct examples of pairs $(C,E)$ as in Theorem 1. In this section we explain how a computer calculation gives our second step toward the proof, by greatly reducing the number of possibilities.
Proposition 12.
Let $C$ and $E$ be as in Theorem 1, let $\varphi\colon C\to E$ be a degree-$2$ map, let $F$ be as in Proposition 3, and let $\Delta_{E}$ and $\Delta_{F}$ be the discriminants of the endomorphism rings of $E$ and $F$, respectively. Then the pair $(\Delta_{E},\Delta_{F})$ is one of the following:
$(-3,-3)$ $(-7,-7)$ $(-8,-72)$ $(-12,-12)$ $(-16,-16)$ $(-20,-20)$
$(-4,-4)$ $(-8,-8)$ $(-11,-11)$ $(-12,-48)$ $(-16,-64)$ $(-24,-24)$
$(-4,-100)$ $(-8,-32)$ $(-12,-3)$ $(-16,-4)$ $(-19,-19)$ $(-36,-36).$
In the cases where $\operatorname{End}E$ and $\operatorname{End}F$ have the same discriminant $\Delta$, then $E$ and $F$ are isomorphic to one another when $\Delta\in\{3,4,7,8,11,12,16,19,20\}$, and $E$ and $F$ are not isomorphic to one another when $\Delta\in\{24,36\}$.
Proof.
Our proof is computational, and Magma programs for carrying out the computation are available at https://github.com/everetthowe/many-maps.
We narrow down our possibilities by using a weakened form of Corollary 4. Given two elliptic curves $E$ and $F$, if there exists an isomorphism $\psi\colon E[2]\to F[2]$ such that $E$, $F$, and $\psi$ are as in Proposition 3, then for every $n>1$ there exists an endomorphism $\alpha_{n}$ of $E$ and a homomorphism $\beta_{n}\colon E\to F$ such that $2n=\deg\alpha_{n}+\deg\beta_{n}$, and such that (6) holds. It follows that (7) also holds, and this is the weaker condition that we will check.
Given the $j$-invariants of $E$ and $F$, we can use the classical modular polynomials $\Psi_{m}$ to determine whether there are endomorphism of $E$, and isogenies $E\to F$, of any given (small) degree and with kernels containing a given number of $2$-torsion points. There is a cyclic isogeny of degree $m$ between two curves with $j$-invariants $j_{1}$ and $j_{2}$ if and only if $\Psi_{m}(j_{1},j_{2})=0$, and every noncyclic isogeny can be factored into a cyclic isogeny composed with multiplication by a rational integer, and the parity of the degree of the cyclic isogeny, and the parity of the rational integer, tell us the size of the $2$-torsion of the kernel.
To prove the proposition, we run through all $(\Delta,p)$ pairs listed in Proposition 6. For each $\Delta$, we construct the number field that contains the $j$-invariants of the elliptic curves $E$ whose endomorphism rings have discriminant $\Delta$; this is simply the number field define by the Hilbert class polynomial for $\Delta$, whose roots are precisely the $j$-invariants in question. Then, for one such root $j_{E}$, we use the classical modular polynomial $\Psi_{p}$ to find the $j$-invariants $j_{F}$ of the elliptic curves $F$ that are $p$-isogenous to $E$.
For each such pair $(j_{E},j_{F})$, we use the method sketched above to compute the set $S_{E}$ of all pairs $(m,d)$ of integers such that there is an endomorphism of $E$ of degree $m$ and with kernel containing exactly $d$ points of order $2$, for $m\leq 62$. We compute the analogous set $S_{F}$ corresponding to morphisms $E\to F$. Then, for every $n$ from $2$ to $31$, we check to see whether we can find an $(m_{1},d_{1})\in S_{E}$ and an $(m_{2},d_{2})\in S_{F}$ with $d_{1}=d_{2}$ and with $m_{1}+m_{2}=2n$.
For all of the pairs $(j_{E},j_{F})$ that meet this requirement, we output the discriminants of the endomorphism rings of $E$ and $F$, and we note whether $j_{E}=j_{F}$. (We can compute the discriminant of $\operatorname{End}F$ by finding the discriminant whose Hilbert class polynomial is equal to the minimal polynomial of $j_{F}$.) The computation gives us the list of discriminant pairs listed in the proposition, and tells us whether $E\cong F$ when the discriminants are equal. ∎
Remark 13.
Magma includes many of the classical modular polynomials $\Psi_{m}$ for $m<62$ in its standard distribution, but not all of them. It does include those for prime powers less than $61$. We obtained $\Psi_{61}$ from Andrew Sutherland’s web page; it was calculated using the methods of [6]. For $m$ with more than one prime factor, we write $m=ab$ for coprime $a$ and $b$, and note that $\Psi(x,y)$ can be computed by taking the $z$-resultant of $\Psi(x,z)$ and $\Psi(y,z)$.
All of the discriminants listed in Proposition 12 have class number at most $2$, and in fact the discriminants on the list that have class number $2$ are all fundamental discriminants. This makes it a simple matter to find the elements of the fundamental domains ${\mathcal{F}}_{1}$ and ${\mathcal{F}}_{2}$ that correspond to elliptic curves with one of these endomorphism rings. Namely, if an order ${\mathcal{O}}={\mathbf{Z}}[\theta]$ has class number $1$, then the lattice $\langle 1,\theta\rangle$ has CM by $\mathcal{O}$}, and the image $\vartheta$ of $\theta$ in the upper half-plane gives rise to the unique elliptic curve over $\mathbf{C}$} with CM by $\mathcal{O}$}. It is a simple matter to find the element $\tau$ of ${\mathcal{F}}_{1}$ that is in the $\operatorname{PSL}_{2}({\mathbf{Z}})$-orbit of $\vartheta$. The elements of ${\mathcal{F}}_{2}$ that correspond to the elliptic curve with CM by $\mathcal{O}$} are simply the images of $\tau$ under the Möbius transformations listed in Figure 2.
If ${\mathcal{O}}={\mathbf{Z}}[\theta]$ is a maximal order with class number $2$, then in addition to the elements given above, we also have values of $\tau$ in the upper half-plane that correspond to the non-principal ideals of $\mathcal{O}$}. Given a nonprincipal prime ideal $I$ of $\mathcal{O}$} lying over a rational prime $p$, we write $I=\langle p,\gamma\rangle$ for some $\gamma\in{\mathcal{O}}$. Then we let $\vartheta$ be the image of $\gamma/p$ in the upper half-plane, and in the same way as above find the elements $\tau$ of ${\mathcal{F}}_{1}$ and ${\mathcal{F}}_{2}$ that lie in the $\operatorname{PSL}_{2}({\mathbf{Z}})$-orbit of $\vartheta$.
We are now in a position to compute the values of $\tau$ and $\sigma$ such that the period lattice (2) with polarization (3) corresponds to the Jacobian of a curve with maps of every degree $n>1$ to the elliptic curve corresponding to the element $\tau$ of ${\mathcal{F}}_{1}$.
Proposition 14.
The pairs $(\tau,\sigma)$ listed in Table 1 are exactly the elements of ${\mathcal{F}}_{1}\times{\mathcal{F}}_{2}$ such that the period matrix (2) and polarization (3) correspond to a curve from Theorem 1 whose associated elliptic curve corresponds to $\tau\in{\mathcal{F}}_{1}$.
Proof.
Our proof is computational, and Magma programs for carrying out the computation are available at https://github.com/everetthowe/many-maps.
For every pair $(\Delta_{E},\Delta_{F})$ in Proposition 12, we let $K={\mathbf{Q}}(\sqrt{\Delta_{E}})\cong{\mathbf{Q}}(\sqrt{\Delta_{F}})$ and we specify an embedding $K$ into $\mathbf{C}$} by choosing one of the square roots of $\Delta_{E}$ in $K$ and declaring that it has positive imaginary part. Then we compute the values of $\tau$ in $K\cap{\mathcal{F}}_{1}$ corresponding to $\Delta_{E}$ as described above, and the values of $\rho\in K\cap{\mathcal{F}}_{1}$ corresponding to $\Delta_{F}$. For each $\rho$ we compute its images $\sigma$ in $K\cap{\mathcal{F}}_{2}$; but if $\Delta_{E}=\Delta_{F}$ and $\Delta_{E}\in\{3,4,7,8,11,12,16,19,20\}$ we only do so when $\rho=\tau$, and if $\Delta_{E}\neq\Delta_{F}$ and $\Delta_{E}\in\{24,36\}$ we only do so when $\rho\neq\tau$.
For each pair $(\tau,\sigma)$ we obtain, we perform the following calculation. Let $\Lambda$ be the lattice in $K^{2}\subset{\mathbf{C}}^{2}$ generated by the vectors
$b_{1}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(1,0),\qquad b_{2}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(0,1),\qquad b_{3}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(\tau/2,1/2),\qquad b_{4}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(1/2,\sigma/2).$
We write elements of $\Lambda$ as $\mathbf{Z}$}-linear combinations of these vectors.
Let $\Lambda_{E}$ be the lattice in $K\subset{\mathbf{C}}$ generated by $1$ and $\tau$. If $\Lambda$, with its principal polarization (3), is the Jacobian of a genus-$2$ curve $C$, then the maps from $C$ to $E$ that take a fixed base point to the origin of $E$ correspond to embeddings of $\Lambda_{E}$ into $\Lambda$. Such an embedding is determined by where it sends $1\in\Lambda_{E}$, and the image $x\in\Lambda$ must have the property that $\tau x\in\Lambda$; in other words, $x$ must be an element of the sublattice ${\mathrm{M}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Lambda\cap\tau^{-1}\Lambda$ of $\Lambda$.
If $\Lambda$ is the Jacobian of a curve $C$, then the degree of the map of $C$ to $E$ corresponding to an element $x$ of $\mathrm{M}$} is equal to the value of the pairing (3) applied to $\tau x$ and $x$. Write $\langle\tau x,x\rangle$ for the value of this pairing. If we compute four elements $c_{1},c_{2},c_{3},c_{4}$ of $\Lambda$ that generate $\mathrm{M}$}, we can then compute the Gram matrix of the quadratic form $q$ on ${\mathbf{Z}}^{4}$ that sends a vector $(n_{1},n_{2},n_{3},n_{4})$ to $\langle\tau x,x\rangle$, where $x=n_{1}c_{1}+\cdots+n_{4}c_{4}$.
Given the Gram matrix of $q$, we can compute all of $v\in{\mathbf{Z}}^{4}$ with $q(v)\leq 31$. Let $S$ be this set of vectors, and let $V$ be the set $\{q(v):x\in S\}$. If $1\in V$ then we know that $\Lambda$ is not the Jacobian of a genus-$2$ curve $C$, because there can be no degree-$1$ map from a genus-$2$ curve to $E$. On the other hand, if $V$ does not contain $1$, then $\Lambda$ does correspond to a curve $C$, and if $V$ does not contain every integer between $2$ and $31$ then $C$ certainly does not have maps of every degree to $E$.
Thus, we can remove from consideration every pair $(\tau,\sigma)$ for which the set $V$ is not equal to $\{2,\ldots,31\}$.
On the other hand, for every pair $(\tau,\sigma)$ for which the set $V$ is equal to $\{2,\ldots,31\}$, we can check to see whether the quadratic form $q$ is isomorphic to one of the forms $q_{1}$, $q_{2}$, $q_{3}$, or $q_{4}$. When we perform this calculation, we find that in fact every such $q$ is isomorphic to one of the forms $q_{i}$. So for each such $q$, we output the values $\Delta_{E}$, $\Delta_{F}$, $\tau$, $\sigma$, and $i$.
When we do so, we find that the output matches Table 1. This proves Theorem 1. ∎
5. Models of the curves
Given values $\Delta_{E}$, $\Delta_{F}$, $\tau$, and $\sigma$ from a row of Table 1, let $K$ be a number field in which the Hilbert class polynomials of $\Delta_{E}$ and $\Delta_{F}$ split. We can compute models over $K$ for elliptic curves $E$ and $F$ with period lattices $\Lambda_{E}$ and $\Lambda_{F}$ homothetic to $\langle 1,\tau\rangle$ and $\langle 1,\sigma\rangle$, respectively. Let $L$ be an extension of $K$ over which the $2$-torsion points of $E$ and $F$ are rational. By complex approximations we can identify the $2$-torsion points $P_{1}$, $P_{\tau}$, and $P_{1+\tau}$ of $E(L)$ corresponding to the values $1/2$, $\tau/2$, and $(1+\tau)/2$ modulo $\Lambda_{E}$, and similarly we can compute the analogously-defined $2$-torsion points $Q_{1}$, $Q_{\sigma}$, and $Q_{1+\sigma}$ on $F(L)$. Let $\psi$ be the isomorphism $E[2]\to F[2]$ that sends $P_{1}$ to $Q_{\sigma}$ and $P_{\tau}$ to $Q_{1}$. Then we can use the formulas from [15, §2] to compute a curve $C$ over $L$ that corresponds to $E$, $F$, and $\psi$ as in Proposition 3.
Once we have a curve $C$ in hand, we can try to find a twist of it that is rational over its field of moduli and that has a relatively simple defining equation. (“Relatively simple” is an inexact expression, so creating these models is not an exact science.)
By this method, we have found the models for the curves in Theorem 1 that we present in Table 2. Since we define the curves in terms of elements of abstract number fields, and not by specific complex numbers, each of the equations corresponds to several curves from Theorem 1 — namely, the ones with the same values of $\Delta_{E}$ and $\Delta_{F}$.
$\Delta_{E}$ $\Delta_{F}$ Polynomial $f$
4-4 100-100 $x^{6}-3x^{4}+(2+r^{24})x^{2}-r^{24}$
where $r^{2}-r-1=0$
8-8 32-32 $x^{6}+(32r^{3}-31r^{2}+8r-18)x^{4}+(8r^{3}-8r^{2}+16r)x^{2}+8$
where $r^{4}-2r^{2}-1=0$
8-8 72-72 $x^{6}+(-2r-33)x^{4}+(-116r-189)x^{2}-2r+5$
where $r^{2}-6=0$
12-12 3-3 $x^{6}+(3r-6)x^{4}+(-12r+9)x^{2}+4$
where $r^{2}+1=0$
16-16 4-4 $x^{6}+(9r+12)x^{4}+(144r-60)x^{2}+64$
where $r^{2}-2=0$
20-20 20-20 $x^{5}+5x^{3}+5x$
24-24 24-24 $x^{6}-21x^{4}+48x^{3}-45x^{2}+48x-23$
36-36 36-36 $x^{5}+(8r-12)x^{4}-(73r+6)x^{3}+(168r+252)x^{2}-(72r+423)x$
where $r^{2}+3=0$
Table 2. For each pair $\Delta_{E}$, $\Delta_{F}$, we give a polynomial $f$ such that $y^{2}=f$ is an equation for the corresponding curves $C$ from Table 1
Example 15.
For the first curve on the list, we will present a basis for the rank-$4$ $\mathbf{Z}$}-module of maps from $C$ to $E$ that take the point $(1,0)$ to the origin of $E$. We give the reasonably simple formulas here; Magma code that verifies (8) can be found in the GitHub repository mentioned in Section 4.
Let $i$ and $s$ satisfy $i^{2}=-1$ and $s^{2}=5$, and let $r=(s+1)/2$, so that $r^{2}-r-1=0$. Our curve $C$ is
$y^{2}=x^{6}-3x^{4}+(2+r^{24})x^{2}-r^{24},$
and we let $E$ be the elliptic curve
$w^{2}=z^{3}+9sz$
with $j$-invariant $1728$.
Define rational function $P_{2}$, $Q_{2}$, $P_{3}$, and $Q_{3}$ by
$P_{2}$ $\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{36sr^{6}}{x^{2}-1}$ $Q_{2}$ $\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{-18sr^{3}}{(x^{2}-1)^{2}}$
$P_{3}$ $\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{-3r(x+1)(x^{2}-6r^{3}x+r^{12})}{(x-1)(sx+r^{6})^{2}}$ $Q_{3}$ $\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{-9r((1-2s)x^{2}-2x-r^{6})}{(x-1)^{2}(sx+r^{6})^{3}}$
and define maps from $C$ to $E$ by
$\varphi_{1}\colon\quad(x,y)$ $\to(\phantom{-}P_{2},\phantom{i}yQ_{2})$
$\varphi_{2}\colon\quad(x,y)$ $\to(-P_{2},iyQ_{2})$
$\varphi_{3}\colon\quad(x,y)$ $\to(\phantom{-}P_{3},iyQ_{3})$
$\varphi_{4}\colon\quad(x,y)$ $\to(-P_{3},\phantom{i}yQ_{3}).$
Then one can check that for integers $a,b,c,d$, we have
(8) $\operatorname{deg}(a\varphi_{1}+b\varphi_{2}+c\varphi_{3}+d\varphi_{4})=2a^{2}+2b^{2}+3c^{2}+3d^{2}+2ad+2bc=q_{2}(a,b,c,d),$
where $q_{2}$ is the quadratic form given in the introduction.
6. Quaternary quadratic forms representing
all integers greater than 11
In this section we will prove Proposition 2. It is easy to check that each of the four forms $q_{1}$, $q_{2}$, $q_{3}$, $q_{4}$ represents the integer $4$, and it is clearly the case that if a form represents $n$ then it also represents $4n$. Thus, it will suffice for us to show that each of the four forms represents every integer $n>1$ that is not a multiple of $4$. We give a separate argument for each of the four forms.
6.1. The quadratic form $q_{1}$
Recall that
$q_{1}=2w^{2}+3x^{2}+3y^{2}+4z^{2}+2xy.$
Suppose $n>1$ is not a multiple of $4$. Let $d$ be the integer defined by
$d=\begin{cases}0&\text{if $n\equiv 2,3,5\bmod 8$}\\ 1&\text{if $n\equiv 1,6,7\bmod 8$.}\end{cases}$
Then $n-4d^{2}$ is a positive integer and $n-4d^{2}\equiv 2,3,5\bmod 8$; a result of Dickson [10, Theorem VI] then shows that we may write
$n-4d^{2}=a^{2}+2b^{2}+2c^{2}$
for some integers $a,b,c$. By considering the right-hand sise of this equality modulo $8$, it is easy to check that $b$ and $c$ cannot both have the opposite parity to $a$, so by switching $b$ and $c$ if necessary we can ensure that $a\equiv b\bmod 2$. Now we simply set
$w$ $=c$ $y$ $=(b-a)/2$
$x$ $=(a+b)/2$ $z$ $=d$
and note that $n=q_{1}(w,x,y,z)$.
6.2. The quadratic form $q_{2}$
Recall that
$q_{2}=2w^{2}+2x^{2}+3y^{2}+3z^{2}+2wz+2xy.$
Suppose $n>1$ is not a multiple of $4$. Let $d$ be the integer defined by
$d=\begin{cases}0&\text{if $3n\equiv 1,2,5,6,7\bmod 8$}\\ 1&\text{if $3n\equiv 3\bmod 8$.}\end{cases}$
Then $3n-5d^{2}$ is a positive integer and $3n-5d^{2}\equiv 1,2,5,6,7\bmod 8$. Another result of Dickson [9, Theorem 15] shows that we may write
$3n-5d^{2}=a^{2}+b^{2}+5c^{2}$
for some integers $a,b,c$. By changing the signs of $a$, $b$, and $c$, if necessary, we may assume that none of $a$, $b$, $c$, $d$ is congruent to $2$ modulo $3$. Note that the preceding equality shows that
$a^{2}+b^{2}\equiv c^{2}+d^{2}\bmod 3,$
so by exchanging $a$ and $b$, if necessary, we may assume that $a\equiv c\bmod 3$ and $b\equiv d\bmod 3$. Now let
$w$ $=c$ $y$ $=(b-d)/3$
$x$ $=d$ $z$ $=(a-c)/3$
and note that $n=q_{2}(w,x,y,z)$.
6.3. The quadratic form $q_{3}$
Recall that
$q_{3}=2w^{2}+3x^{2}+3y^{2}+4z^{2}+2wx+2wy+2xz+2yz.$
Let $n>1$ be an integer that is not a multiple of $4$. Let $d$ be the integer defined by
$d=\begin{cases}0&\text{if $n\equiv 2,3,6,7\bmod 8$}\\ 1&\text{if $n\equiv 1,5\bmod 8$.}\end{cases}$
Then $n-3d^{2}$ is a positive integer and $n-3d^{2}\equiv 2,3,6,7\bmod 8$. Yet another result of Dickson [10, Theorem XI] shows that we may write
$n-3d^{2}=a^{2}+2(b^{2}+bc+c^{2})$
for some integers $a,b,c$. Now, $b$ and $c$ cannot both be even, because in that case we would have $n-3d^{2}\equiv a^{2}\bmod 8$, while we know that $n-3d^{2}$ is not congruent to a square modulo $8$. By replacing $(b,c)$ with (b+c,c)(b+c,-c) if necessary, we can ensure that one of the numbers $b,c$ is even and the other odd. Then by switching $b$ and $c$, if necessary, we can ensure that $a+c+d$ is even. Now set
$w$ $=b+(a+c+d)/2$ $y$ $=c-(a+c+d)/2$
$x$ $=-(a+c+d)/2$ $z$ $=d$
and note that $n=q_{2}(w,x,y,z)$.
6.4. The quadratic form $q_{4}$
Recall that
$q_{4}=2w^{2}+3x^{2}+4y^{2}+6z^{2}-2wx+2wz+2xy+4yz.$
Our proof in this case depends on the parity of $n$. First let us suppose that $n>1$ is an even integer that is not a multiple of $4$. Then $n$ is congruent to $2$ or $6$ modulo $8$, so Legendre’s three-square theorem [27, p. 378] says that we may write
$n=a^{2}+b^{2}+c^{2}$
for some integers $a,b,c$. By permuting these integers and changing their signs, if necessary, we may assume that $a\equiv b\bmod 3$. Then we may set
$w$ $=(2a+b)/3$ $y$ $=(a-b+3c)/6$
$x$ $=0$ $z$ $=(b-a)/3.$
These numbers are integers; the only thing that may not be clear immediately is whether $y$ is integral at $2$, but that can be verified by noting that
$a-b+3c\equiv a+b+c\equiv a^{2}+b^{2}+c^{2}\equiv n\equiv 0\bmod 2.$
One can easily check that $n=q_{4}(w,x,y,z).$
Now suppose that $n>1$ is odd. Let $d=3$, and note that the three-square theorem shows that we may write
$4n-d^{2}=a^{2}+b^{2}+c^{2}$
for some integers $a,b,c$. Considering this equality modulo $4$, we see that $a$, $b$, and $c$ must all be odd, and by permuting them and changing their signs (if necessary) we can assume that $a\equiv b\bmod 3$ and $a\equiv b+c+d\bmod 4$. Now set
$w$ $=(2a+b+d)/6$ $y$ $=(a-b+3c-d)/12$
$x$ $=d/3$ $z$ $=(b-a)/6.$
Our assumptions on $a$, $b$, $c$, and $d$ show that $w,x,y,z$ are integers, and it is easy to check that $n=q_{4}(w,x,y,z).$
This proves Proposition 2.∎